Suppose the width of a certain rectangle is 1 inch more than one-fourth of its lenght. the perimeter of the rectangle is 42 inches. find the lenght and width of the rectangle.
Answer by checkley77(12844)

LET WIDTH=L/4+1 WHERE L IS THE LENGTH.
2W+2L=42
2(L/4+1)+2L=42
2L/4+2+2L=42
2L/4+2L=42-2
(2L+4*2L)/4=40
(2L+8L)/4=40
10L/4=40
10L=4*40
10L=160
L=160/10
L=16 FOR THE LENGTH.
W=16/4+1
W=4+1
L=5 FOR THE LENGTH.
PROOF:
2*16+2*5=42
32+10=42
42=42